### v1

# Solving a Stackelberg Game on Transportation Networks in a Dynamic Crime Scenario: A Mixed Approach on Multi-Layer Networks

Abstract
Interdicting a criminal with limited police resources is a challenging task as the criminal changes location over time. The size of the large transportation network further adds to the difficulty of this scenario. To tackle this issue, we consider the concept of a layered graph. At each time stamp, we create a copy of the entire transportation network to track the possible movements of both players, the attacker and the defenders. We consider a Stackelberg game in a dynamic crime scenario where the attacker changes location over time while the defenders attempt to interdict the attacker on his escape route. Given a set of defender strategies, the optimal attacker strategy is determined by applying Dijkstra’s algorithm on the layered networks. Here, the attacker aims to minimize while the defenders aim to maximize the probability of interdiction. We develop an approximation algorithm on the layered networks to find near-optimal strategy for defenders. The efficacy of the developed approach is compared with the adopted MILP approach. We compare the results in terms of computational time and solution quality. The quality of the results demonstrates the need for the developed approach, as it effectively solves the complex problem within a short amount of time.
keywords:
Resource allocation, Multi-layer time expanded network, Vehicle routing and scheduling, Network optimization, Exact algorithm, Approximation algorithm
1 Introduction
We consider a Stackelberg game where the defender is the leader, and the attacker is the follower. The assumption is that the defenders have information only about the crime location in a large transportation network. In this Stackelberg game, first, the defenders commit to a strategy, and then the attacker generates the best response to the given defender strategy. We use the concept of a multi-layer network in which, at each time-stamp, we create a copy of the entire network. We consider edge length as the time factor to create connections between these multi-layer networks. This means that depending on the edge length, we choose the layers from which the start and end nodes of that particular edge are selected. In this way, for all edges in the original transportation network, we create corresponding edges in the multi-layer network. We consider a set of strategies for both players, i.e., the defenders and the attacker. For the defender, we use a mixed strategy set where each strategy is assigned a mixed probability, with the sum of these probabilities equaling one. For the attacker, we consider a pure strategy set.
To generate the optimal attacker strategy, we first assign weights to the nodes depending on the mixed probabilities and the nodes’ presence in the defender strategies. Considering all defender strategies, we sum up the probabilities if any node is present in multiple defender strategies. Then, we assign the weight of each node to all of its incoming edges. Finally, we apply Dijkstra’s algorithm on this network to find the optimal strategy for the attacker. For the defender, we create a multi-layer network in the same way as for the attacker. Weights are assigned to all nodes based on the attacker strategies, with each attacker strategy given equal probability. We develop a novel approximation algorithm for the defender to generate a near-optimal defender strategy. Thus, for the defender, the problem can be considered as finding a near-optimal path with the aim of covering at least one node from each attacker strategy. For this, we use different colors to represent each attacker strategy. The developed defender strategy includes as many different colored vertices as possible. We consider a Stackelberg game instead of a zero-sum game. The advantage of the Stackelberg game over the zero-sum game is that it allows for a more realistic representation of real-world scenarios where one player has more information, resources, or power than the other player. In this Stackelberg game, the defenders have an advantage because they can consider the attacker’s potential moves and responses when making their decisions. This can result in a more efficient outcome than in a zero-sum game, where the players are equally matched.
Due to the complex transportation network and limited police resources, generating efficient strategies for both players is a challenging task. We can summarize the main contributions of this paper as follows: (1) To the best of our knowledge, this is the first attempt to solve a Stackelberg game considering the escape interdiction game (EIG) problem using the multi-layer network concept, with a focus on developing solution methodologies. (2) We develop an optimal approach for the attacker and a near-optimal approach for the defenders, and we are not aware of any previous implementations for escape interdiction problems on multi-layer network. (3) We generate a multi-layer network to create efficient strategies for both players. (4) We use grids dataset for the case study and show that our MLN-EIGS approach outperforms MILP-EIGS approach in terms of solution quality, with a minimal performance gap.
The paper is organized as follows. In Section 2, we present the relevant research. In Section 3, we define the problem description and modeling. In Section 4, we describe the solution methodology. Section 5 presents the quality of the results. We conclude this research in Section 6.
2 Related Work
Security games play an important role in providing social security. Zhang et al. [3] consider a zero-sum game for the escape interdiction problem and develop a MILP-based solution approach to provide an optimal solution. To demonstrate the efficacy of their methodology, they generate optimal solutions on grids of different sizes. Samanta et al. [1] consider the same zero-sum game problem and develop a meta-heuristic-based solution approach to provide a scalable near-optimal solution in a time-efficient manner. For the same escape interdiction problem, Samanta et al. [2] develop a simulation-based approach to generate a scalable solution to increase security in a large transportation network.
Recent literature focuses on Stackelberg games considering the escape interdiction problem to decrease the crime rate in society. For example, Zychowski and Mandziuk [4] consider search games (SEG) on directed graphs. They consider multiple defender resources and one attacker, where the attacker aims to reach one of several predefined target points from a fixed location. They develop a genetic algorithm-based heuristic approach to provide a near-optimal solution on synthetic datasets.
Basilico et al. [9] consider a leader-follower game and formulate the problem as a mathematical programming model. They use optimization software tools to solve the model by generating the optimal strategies for both leader and follower. Letchford and Conitzer [11] consider security games on graphs and develop a polynomial-time algorithm to generate optimal strategies for players. Similarly, Iwashita et al. [10] consider security games on graphs and develop an algorithm to reduce the graph by eliminating unnecessary edges, providing a time-efficient, scalable near-optimal solution. In the same vein, Karwowski and Mandziuk [18] consider an evader-defender Stackelberg game model and develop a Monte Carlo Tree Search approach to provide efficient patrolling schemes.
Tsai et al. [13] consider an attacker-defender Stackelberg game model and develop a linear program to generate optimal mixed strategies for players by allocating limited resources optimally, with a case study on part of the Mumbai road network. Similarly, Shieh et al. [12] develop a game-theoretic system to provide security with limited resources in the port of Boston and also test its efficacy in the port of New York. They schedule patrols efficiently to provide optimal mixed strategies for players, considering an attacker-defender Stackelberg game model. Similar papers focus on Stackelberg security games to provide security in society (e.g., Cermak et al. [16], Lou et al. [19], Sinha et al. [14], ˇCern‘y et al. [17], Zhang and Malacaria [15]).
Conitzer and Sandholm [5] show that exact approaches are not suitable because of the NP-hard nature of this problem, as it faces scalability issues due to the high time complexity. Paruchuri et al. [6] develop a MILP-based exact algorithm to solve Bayesian Stackelberg security games. Bosansky and Cermak [7] also consider Stackelberg games and develop algorithms based on LPs and MILP to generate Strong Stackelberg Equilibrium (SSE) and perform a case study at Los Angeles International Airport, considering security scheduling.
The layered graph concept is a useful tool for solving complex problems on transportation networks in a time-efficient manner. For example, Saito et al.[8] consider the problem of selecting important nodes in a network and construct a layered graph from the original graph, where each layer is added on top as time proceeds to demystify the complex problem. Hence, we focus on solving a Stackelberg escape interdiction game using the layered graph concept.
3 Problem Description
We consider a two-player Stackelberg game where this sequential game is played between multiple defenders $(D)$ and one attacker $(A)$. The defenders have a finite set of actions $\acute{S}$, and the attacker has a finite set of actions represented by $\acute{A}$. $S$ and $A$ represent the sets of probability distributions over the defender and attacker actions, respectively. The game is modeled as a general-sum game, where $x\in S$ and $y\in A$ are the probability distributions over $\acute{S}$ and $\acute{A}$, respectively.
For $a_{j}=(v_{j},t^{a}_{j})$ and $s^{r}_{i}=(v^{r}_{i},t^{r,in}_{i},t^{r,out}_{i})$, $d_{r}$ reaches the attacker at node $v^{r}_{i}$ if $v^{r}_{i}$ = $v_{j}$ and $t^{r,in}_{i}\leq t^{a}_{j}\leq t^{r,out}_{i}$. The defender gets -1 if the attacker successfully escapes from the transportation network; otherwise, the defender’s utility value is 0 (see Eq. 1). Considering this case, we define $Z_{s^{r}_{i},a_{j}}=1$, otherwise $Z_{s^{r}_{i},a_{j}}=0$.
Here, $y$ and $A$ are used to calculate the expected utility of the defender, which is represented in Eqs. (2) and (3).
$U_{d}\;(S,A)=\begin{cases}0,&\text{if $\exists$$Z_{s^{r}_{i},a_{j}}$ = 1, $a_{j}$ $\in$ A, $s^{r}_{i}$ $\in$ $S^{r}$, r $\in$ R}.\\ -1,&\text{ otherwise}.\end{cases}$ (1)
$U_{d}\;(x,A)=\sum\limits_{S\in\textbf{S}}U_{d}(S,A)\;x_{s}$ (2)
$U_{d}\;(x,y)=\sum\limits_{A\in\textbf{A}}U_{d}(x,A)\;y_{A}$ (3)
Therefore, we need an optimization model so that the defenders can find their best optimal strategy by solving the following linear problem (see Eqs. 4-6).
$max\qquad U^{*}$ (4)
$s.t.\qquad U^{*}\leq U_{d}\;(x,A)\qquad\forall A\in A^{{}^{\prime}}$ (5)
$\sum\limits_{S\in\textbf{S}}x_{s}=1,x_{s}\geq 0\qquad\forall S\in S^{{}^{\prime}}$ (6)
The expected utility of the defender and attacker are $U_{d}$ and $U_{a}$, respectively. For any pair of strategies $(x,y)$, the expected payoff of the defender is denoted by $U_{d}(x,y)$.
First, in Eq. (7), the defender commits to a strategy $x\in X$, and given such an $x$, the attacker chooses his strategy from the best-response set $BR(x)$ where
$BR(x)=\;argmax_{y\in Y}\;U_{a}\;(x,y)$ (7)
Second, in Eq. (8), to maximize utility, the defender chooses the strategy $x$ (the best response of the defender) to the attacker’s best response.
$max_{x\in S}\;U_{d}\;(x,y)\qquad s.t.\;y\in BR(x)$ (8)
Third, in a Strong Stackelberg Equilibrium (SSE), we assume that the follower (attacker) breaks ties in favor of the leader (defender) (see Eq. 9). In that case, the optimization problem is
$max_{x\in S,\;y\in BR(x)}\;U_{d}\;(x,y)$ (9)
SSE is popular because it is always guaranteed to exist ((Conitzer [20]).
4 Proposed Mixed Approach on Multi-Layer Networks
This section presents the proposed methodology for the considered escape interdiction problem which is formulated as a Stackelberg game model. A novel mixed approach on layered graph concept is developed to solve the Stackelberg model, where the defender is the leader and the attacker is the follower. For the defender, the developed approximation algorithm finds the efficient movement of the defender from a random initial location. For the attacker, the initial strategy is any random path from the crime location to any exit point. An exact algorithm is developed to generate the optimal path for the attacker using Dijkstra’s algorithm on the time-expanded network. The new defender strategy is generated using a greedy algorithm on the time-expanded network. The defender’s mixed strategy set is updated based on the recalculated utility. The defender then selects the best strategy to commit to. This best defender strategy is given as input to the attacker. The attacker tries to generate a new optimal strategy using the exact approach, taking the new defender strategy as input. The attacker updates the pure attacker strategy set, where all strategies are of equal probability. The algorithm checks if, for $K$ consecutive iterations, the game utility is below a threshold value named $\epsilon$. The value of $\epsilon$ is set to 0.05 for our experiments. If this convergence criterion is met, the algorithm converges, and we return the best defender strategy with the probability of interdiction and total runtime (see Algorithm 1). The flowchart is shown in Fig. 1.
Figure 1: Flowchart of Stackelberg Game Model
Input: Initialize the initial strategy sets of defender and attacker $S^{\prime},A^{\prime}$;
Output: $U_{d}:$ Defenders’ game utility;
repeat
       $(x,y)\leftarrow CoreLP(S^{\prime},A^{\prime})$;
       Defender commits to the best strategy $x^{\star}\in x$ which provides maximum utility of defender ;
       $BR:\;A^{*}\leftarrow{ExactAO(x^{\star})}$;
       if $A^{*}\neq\emptyset$ then
             $A{{}^{\prime}}\leftarrow A^{{}^{\prime}}\cup A^{*}$;
       end if
      $S^{*}\leftarrow{ApproxDO(A^{{}^{\prime}})}$;
       if $S^{*}\neq\emptyset$ then
             $S^{{}^{\prime}}\leftarrow S^{{}^{\prime}}\cup S^{*}$;
       end if
until convergence;
return $U_{d}$.
Algorithm 1 MLN-EIGS Algorithm
To create a benchmark, we construct a novel Stackelberg game where the defender is the leader, and the attacker is the follower. The best strategies for the attacker and defender are generated using bestAO and bestDo, which are the MILP approaches for the attacker and defender, respectively (see Algorithm 2). Zhang et al. [3] develop these optimal approaches to provide optimal strategies for attackers and defenders given a strategy set for each player. Similar to the MLN-EIGS, in this MILP-EIGS the defender commits to the best mixed strategy. The attacker takes this best defender mixed strategy as input, generates the optimal attacker strategy, and updates the pure attacker strategy set. Then, we calculate the game utility, and the game continues until the game utility is below $\epsilon$ for $K$ consecutive iterations. Here $\epsilon$ acts as a threshold with a value of 0.05. We return the defender utility value once the convergence is reached.
Input: Initialize the initial strategy sets of defender and attacker $S^{\prime},A^{\prime}$;
Output: $U_{d}:$ Defenders’ game utility;
repeat
       $(x,y)\leftarrow CoreLP(S^{\prime},A^{\prime})$;
       Defender commits to the best strategy $x^{\star}\in x$ which provides maximum utility of defender ;
       $BR:\;A^{*}\leftarrow{BestAO(x^{\star})}$;
       if $A^{*}\neq\emptyset$ then
             $A{{}^{\prime}}\leftarrow A^{{}^{\prime}}\cup A^{*}$;
       end if
      $S^{*}\leftarrow{BestDO(A^{{}^{\prime}})}$;
       if $S^{*}\neq\emptyset$ then
             $S^{{}^{\prime}}\leftarrow S^{{}^{\prime}}\cup S^{*}$;
       end if
until convergence;
return $U_{d}$.
Algorithm 2 MILP-EIGS Benchmarking Algorithm
4.1 Efficient Attacker Strategy Design Using Exact Approach on Time Expanded Network
The optimal attacker strategy is generated using the exact approach for the attacker (see Algorithm 3). For each time-stamp, we create a copy of the entire network. Then, we establish connections in this multi-layer network using the edge lengths in the original graph. This algorithm describes how we assign the probability of interdiction to each node. Finally, we apply Dijkstra’s algorithm on this multi-layer network to generate the optimal strategy for the attacker.
Input: Crime node is the “START” node, all exit nodes are the “GOAL” nodes, Original graph $(G_{0})$;
Output: Best attacker strategy having minimum probability of interdiction;
Construction of layered graphs:
for $(i=0;i<t_{max};i++)$ do
       Generate one copy of the original graph/network, labelled as $G_{i+1}$ ;
end for
Connect the Layered Graphs depending on time/distance between adjacent nodes in the original graph:
for all edges in the original graph do
       $L$ = current edge length, $S$ = From node, $T$ = To node (in original graph) ;
       for $(j=0;j<t_{max}-L;j++)$ do
             Select graph $G_{j}$ and graph $G_{j+L}$ ;
             Create edge from node $S$ of $G_{j}$ to node $T$ of $G_{j+L}$ ;
       end for
end for
Update the minimum Prob. of. Interdiction $(1-P)$ value to the node depending on its presence in the defender strategy:
Initial probability of interdiction $(P)$ of all nodes in all layered graphs = 0 ;
for all $N$ defender strategies in the best strategy set do
       for all nodes present in the current strategy do
             $Node=n,t_{in}=In\;Time,t_{out}=Out\;Time$ while $t_{in}<t_{out}$ do
                   Select node $n$ of graph $G_{t_{i}n}$ Update the minimum Prob. of Interdiction of that node $n:(P=P+P_{mix})$ where $P_{mix}$ is the mixed prob. of the current defender strategy. $t_{in}=t_{in}+1$
             end while
       end for
end for
Algorithm 3 Optimal Attacker Strategy Design Using Time Expanded Network
Update the minimum Prob. of. Interdiction $(1-P)$ value to the node depending on its presence in the defender strategy:
for all nodes present in the layered graph do
       if Prob. of interdiction of a node $n$ is $P$ then
             Assign the weight $P$ to all in-coming edges of that node.
       end if
end for
Apply Dijkstra’s algorithm on the developed Multi-Layer Network.
return $U_{d}$.
4.2 Efficient Defender Strategy Design Using Approximation Algorithm on Time Expanded Network
We develop a novel approximation algorithm to generate a near-optimal strategy for the defender in a time-efficient manner. We describe the approach in the following steps. Since the attacker’s strategy set consists of pure strategies, we assign different colors to each of these attacker strategies. Then, we attempt to construct a path for the defender that includes at least one colored vertex to interdict each attacker strategy. This implies that the defender’s strategy contains as many different colored vertices as possible.
  • 1.
    Input
    • (a)
      A directed acyclic weighted graph, where some vertices are colored.
    • (b)
      There are $k$ different colors $(c_{1},c_{2},...,c_{k})$.
    • (c)
      A threshold value of the path length $t$.
    Here, each colored vertex corresponds to a particular attacker’s strategy (a path).
  • 2.
    Goal:
    To find a path whose length is at most $t$ and contains as many different colored vertices as possible.
  • 3.
    Guess:
    This problem is difficult as Dijkstra cannot solve its decision version, which is $NP$-complete. We provide a formal proof in this paper.
We can use the steps below for the approximate algorithm.
  • 1.
    Step1: Find a shortest path to any one of colored vertices from the start vertex (we can use Dijkstra). Assume the path is to vertex $v_{1}$ with color $c_{i_{1}}$.
  • 2.
    Step2: Find a shortest path to any one colored vertices except $c_{i_{1}}$ from $v_{1}$. Assume the path is to vertex $v_{2}$ with color $c_{i_{2}}$.
  • 3.
    Step3: Find a shortest path to any one colored vertices except $c_{i_{1}}$ and $c_{i_{2}}$ from $v_{2}$, and so on, until all colors are visited or the total path length reaches $t$.
If we cannot find a path that covers all colors, we can create a tree (or a set of paths) instead, which corresponds to a non-deterministic strategy for a defender.
In a simplest form, if the path we obtained in the previous method does not cover a subset of colors, we create another path from the initial vertex, which tries to cover these remaining colors only. We repeat this procedure until all colors are covered. Then, the defender flips a coin and chooses one path.
4.3 Theorem: Problem ”color covering” is $NP$-complete
Problem ”color covering”
  • 1.
    Input: A directed graph and the initial vertex. Some vertices are colored. There are $m$ different colors. It is possible that one vertex has multiple colors.
  • 2.
    Output: ”yes” when there exists a path from the initial vertex with length $n$, such that all $m$ colors appear at least one of a vertex on the path. ”no” otherwise.
Theorem: Problem ”color covering” is NP-complete
  • 1.
    Proof idea: Reduction from 3-SAT (which is known to be NP-complete).
Problem 3-SAT
  • 1.
    Input: $n$ boolean variables $(x_{1},\ldots,x_{n})$, $m$ clauses. Each clause is a disjunction of three literals. Each literal is a variable or its negation.
  • 2.
    Output: ”yes” if there exists an assignment of variables that makes all clauses true. ”no” otherwise.
For a given $3-SAT$ instance, we create an $n+1$ level network.
  • 1.
    There is one level-00 vertex, which is the initial vertex.
  • 2.
    There are two level-$i$ vertices (for $i>0$). One vertex corresponds to making variable $x_{i}$ true. The other vertex corresponds to making variable $x_{i}$ false.
  • 3.
    There exists a directed edge from each of level-$i$ vertex to each of level-$(i+1)$ vertices.
  • 4.
    Each clause has its own color.
  • 5.
    If a clause with color $c$ contains $x_{i}$, the ”true” vertex for $x_{i}$ has color $c$.
  • 6.
    If a clause with color $c$ does not contain $x_{i}$, the ”false” vertex for $x_{i}$ has color $c$.
Thus, the $3-SAT$ instance is satisfiable iff there exists a path from the initial vertex with length $n$, which covers all colors.
5 Results and Discussion
In this section, we present the results of the developed approaches. The proposed algorithms are coded in Python 3.6 and tested on a computer equipped with an Intel(R) Core(TM) 3.20 GHz processor and 8 GB RAM, operating under the LINUX environment. All MILPs are solved using CPLEX (version 12.8).
In Fig. 2, we consider a sample network and use Dijkstra’s algorithm on the time-expanded network to generate the optimal attacker strategy. Here, we provide two mixed defender strategies as input with probabilities of 0.6 and 0.4. We demonstrate that our developed exact approach for the attacker can generate the optimal attacker strategy, enabling the attacker to escape without interdiction in a concise amount of time (see Table 1).
Figure 2: Optimal Attacker Strategy Design Using Dijkstra Algorithm on Time Expanded Network.
Table 1:
Optimal Attacker Strategy Design Using Dijkstra Algorithm on Time Expanded Network.
Game Parameters
Network Size: Crime Node: $0\_1$, Police Station: $0\_6$, $T_{max}$: 6, Exit Point: 5
Test Case Optimal Attacker strategy Utility of the Final Optimal Attacker Strategy Run Time (Sec)
1 [$0\_1,3\_3,5\_4$] 0.0 0.004
2 [$0\_1,3\_3,5\_4$] 0.0 0.004
3 [$0\_1,3\_3,5\_4$] 0.0 0.0039
4 [$0\_1,3\_3,5\_4$] 0.0 0.004
5 [$0\_1,3\_3,5\_4$] 0.0 0.004
6 [$0\_1,3\_3,5\_4$] 0.0 0.004
7 [$0\_1,3\_3,5\_4$] 0.0 0.004
8 [$0\_1,3\_3,5\_4$] 0.0 0.004
9 [$0\_1,3\_3,5\_4$] 0.0 0.004
10 [$0\_1,3\_3,5\_4$] 0.0 0.004
In Fig. 3, we consider a sample network and use the approximation algorithm on the time-expanded network to generate the near-optimal defender strategy. Here, we provide three pure attacker strategies as input with equal probability. We demonstrate that our developed approach for the defender can generate an efficient defender strategy that interdicts all attacker strategies quickly (see Table 2).
Figure 3: Defender Strategy Design Using Approximation algorithm.
Table 2:
Defender Strategy Design Using Approximation algorithm.
Game Parameters
Network Size: Crime Node: $0\_1$, Police Station: $0\_6$, $T_{max}$: 6, Exit Point: 5
Test Case Final Defender strategy No. of Non-interdicted Attacker Strategies by the Final Defender Strategy Run Time (Sec)
1 [$0\_6,2\_3,4\_4,5\_5,6\_5$] 0.0 0.0086
2 [$0\_6,2\_3,4\_5,5\_5,6\_5$] 0.0 0.0090
3 [$0\_6,2\_3,4\_5,5\_5,6\_5$] 0.0 0.0081
4 [$0\_6,2\_3,4\_5,5\_5,6\_5$] 0.0 0.0081
5 [$0\_6,2\_3,4\_5,5\_5,6\_5$] 0.0 0.0082
6 [$0\_6,2\_3,4\_4,5\_5,6\_5$] 0.0 0.0080
7 [$0\_6,2\_3,4\_4,5\_5,6\_5$] 0.0 0.0082
8 [$0\_6,2\_3,4\_4,5\_5,6\_5$] 0.0 0.0081
9 [$0\_6,2\_3,4\_4,5\_5,6\_5$] 0.0 0.0083
10 [$0\_6,2\_3,4\_5,5\_5,6\_5$] 0.0 0.0081
We compare the final defenders’ utility generated by the developed MLN-EIGS algorithm with the final defenders’ utility using MILP-EIGS algorithm, which uses the exact approaches named bestDo and bestAo, developed by Zhang et al. [3]. To establish a benchmark, we solve the Stackelberg Game using the bestDo and bestAo exact approaches. We compare both the solution quality and computational time.
We have generated 10 test cases for each grid from 3*3 to 9*9 (see Tables 3, 4, 5, 6, 7, 8, and 9). All test case instances are generated by varying the crime node location, initial location of defenders, $t_{max}$ (maximum run time), and exit node. For each test case, we check the performance gap and the variation in computational time.
Table 3:
Grid: 3*3.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 0.0 0.0 3.71 3.59
2 0.0 0.0 3.69 3.49
3 0.0 0.0 3.72 3.38
4 0.0 0.0 3.78 3.52
5 0.0 0.0 3.72 2.16
6 -1.0 -1.0 3.69 2.54
7 0.0 0.0 3.86 2.23
8 0.0 0.0 3.70 2.11
9 0.0 0.0 3.74 2.13
10 -1.0 -1.0 3.73 2.79
Table 4:
Grid: 4*4.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 -1.0 -1.0 12.54 0.72
2 -1.0 -1.0 9.28 0.71
3 -1.0 -1.0 12.53 0.70
4 -1.0 -1.0 12.19 0.70
5 -1.0 -1.0 12.22 0.52
6 -1.0 -1.0 13.32 0.54
7 -1.0 -1.0 12.93 0.52
8 -1.0 -1.0 11.98 0.51
9 0.0 0.0 12.23 0.55
10 -1.0 -1.0 13.38 0.92
Table 5:
Grid: 5*5.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 0.0 0.0 20.35 7.39
2 0.0 0.0 20.24 9.52
3 0.0 0.0 20.86 7.58
4 0.0 0.0 43.62 2.09
5 0.0 0.0 34.86 2.12
6 -1.0 -1.0 37.86 2.54
7 -1.0 -1.0 34.93 2.48
8 -1.0 -1.0 40.28 1.12
9 -1.0 -1.0 37.98 1.11
10 0.0 0.0 35.12 1.00
Table 6:
Grid: 6*6.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 0.0 0.0 72.75 4.14
2 0.0 -1.0 74.52 4.11
3 -1.0 -1.0 72.39 4.02
4 -1.0 -1.0 72.07 3.97
5 -1.0 -1.0 73.27 2.40
6 0.0 0.0 60.17 5.93
7 0.0 0.0 35.11 1.88
8 0.0 0.0 34.42 4.00
9 -1.0 -1.0 60.49 4.62
10 -1.0 -1.0 34.54 4.61
Table 7:
Grid: 7*7.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 0.0 0.0 231.86 15.51
2 0.0 0.0 234.26 16.09
3 0.0 0.0 233.18 20.97
4 -1.0 -1.0 140.45 31.64
5 0.0 0.0 80.82 14.76
6 0.0 0.0 138.56 15.29
7 0.0 0.0 205.43 15.54
8 0.0 0.0 81.21 15.14
9 -1.0 -1.0 140.92 15.09
10 0.0 0.0 143.20 14.88
Table 8:
Grid: 8*8.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 -1.0 -1.0 437.82 43.97
2 -1.0 -1.0 436.68 38.61
3 0.0 0.0 422.33 30.94
4 0.0 0.0 399.56 37.32
5 -1.0 -1.0 406.86 31.32
6 -1.0 -1.0 305.71 29.39
7 0.0 0.0 260.80 28.47
8 0.0 0.0 440.14 28.58
9 0.0 0.0 288.26 27.74
10 0.0 0.0 420.42 29.52
Table 9:
Grid: 9*9.
Test Case Final Defender Utility Run Time (Sec)
MILP-EIGS MLN-EIGS MILP-EIGS MLN-EIGS
1 0.0 0.0 526.49 48.69
2 -1.0 -1.0 349.04 43.79
3 0.0 0.0 354.68 35.70
4 0.0 -0.5 357.99 42.82
5 0.0 0.0 498.56 44.34
6 0.0 0.0 1069.11 43.36
7 0.0 0.0 735.04 44.10
8 0.0 0.0 795.54 36.17
9 0.0 0.0 846.62 35.70
10 0.0 0.0 577.01 81.17
We show that for these small-size grids, our developed mixed approach named MLN-EIGS can provide an optimal solution like the MILP approach named MILP-EIGS in most cases and takes less computational time compared to MILP-EIGS.
6 Conclusion and Future Directions
In this paper, we consider a Stackelberg game theoretic model to solve security games considering the dynamic crime scenario. This paper proposes a novel optimal routing and resource allocation approach on a transportation network. The optimal strategy for the attacker is generated using an exact approach. To the best of our knowledge, this is the first time an exact approach using the multi-layer network concept is used for the attacker to solve a security game. This developed exact approach consists of Dijkstra’s algorithm and the layered graph concept. The near-optimal strategy for the defender is generated using a novel approximation algorithm and the layered graph concept on a transportation network in a time-efficient manner. The experiments are conducted on 3x3 to 9x9 grids.
The results are compared with the solutions developed using MILP approaches for both attacker and defender in this Stackelberg game scenario. For benchmarking, we formulate the Stackelberg game to provide optimal solutions using the adopted bestAO and bestDO approaches. The contribution of the paper is that the developed MLN-EIGS can offer quality solutions like MILP-EIGS but in a significantly shorter amount of time.
The developed approaches have a few limitations. For example, the proposed approach does not include real-time traffic. Future research can focus on developing efficient algorithms to incorporate traffic into the developed mixed approach. In addition, an exact approach for the defender can also be developed to generate an optimal strategy for the defender on a time-layered network.

### v2

@@
 Abstract
-Interdicting a criminal with limited police resources is a challenging task as the criminal changes location over time.
-The size of the large transportation network further adds to the difficulty of this scenario.
-To tackle this issue, we consider the concept of a layered graph.
-At each time stamp, we create a copy of the entire transportation network to track the possible movements of both players, the attacker and the defenders.
-We consider a Stackelberg game in a dynamic crime scenario where the attacker changes location over time while the defenders attempt to interdict the attacker on his escape route.
+Interdicting a criminal with limited police resources is a challenging task in a dynamic crime scenario, as the criminal changes location over time.
+The vastness of the transportation network adds to the difficulty.
+To address this, we introduce the concept of a layered graph, where at each time step, a duplicate of the transportation network is generated to trace the possible movements of both the criminal/attacker and the police/defenders.
+We model this as a Stackelberg game, where the attacker seeks to escape from the network using one of the predefined exit points, while the defenders attempt to intercept the attacker on his escape route.
 Given a set of defender strategies, the optimal attacker strategy is determined by applying Dijkstra’s algorithm on the layered networks.
-Here, the attacker aims to minimize while the defenders aim to maximize the probability of interdiction.
-We develop an approximation algorithm on the layered networks to find near-optimal strategy for defenders.
+The attacker seeks to minimize, while the defenders aim to maximize the probability of interdiction.
+We propose an approximation algorithm on the layered networks to find near-optimal strategy for defenders.
 The efficacy of the developed approach is compared with the adopted MILP approach.
 We compare the results in terms of computational time and solution quality.
-The quality of the results demonstrates the need for the developed approach, as it effectively solves the complex problem within a short amount of time.
+The quality of the results underscores the necessity of the developed approach, as it efficiently solves this complex problem in a short timeframe.
 keywords:
-Resource allocation, Multi-layer time expanded network, Vehicle routing and scheduling, Network optimization, Exact algorithm, Approximation algorithm
+Resource allocation, Multi-layer time expanded network, Vehicle routing and scheduling, Exact algorithm, Approximation algorithm
 1 Introduction
@@
 The assumption is that the defenders have information only about the crime location in a large transportation network.
-In this Stackelberg game, first, the defenders commit to a strategy, and then the attacker generates the best response to the given defender strategy.
-We use the concept of a multi-layer network in which, at each time-stamp, we create a copy of the entire network.
-We consider edge length as the time factor to create connections between these multi-layer networks.
-This means that depending on the edge length, we choose the layers from which the start and end nodes of that particular edge are selected.
-In this way, for all edges in the original transportation network, we create corresponding edges in the multi-layer network.
-We consider a set of strategies for both players, i.e., the defenders and the attacker.
-For the defender, we use a mixed strategy set where each strategy is assigned a mixed probability, with the sum of these probabilities equaling one.
-For the attacker, we consider a pure strategy set.
-To generate the optimal attacker strategy, we first assign weights to the nodes depending on the mixed probabilities and the nodes’ presence in the defender strategies.
-Considering all defender strategies, we sum up the probabilities if any node is present in multiple defender strategies.
-Then, we assign the weight of each node to all of its incoming edges.
-Finally, we apply Dijkstra’s algorithm on this network to find the optimal strategy for the attacker.
-For the defender, we create a multi-layer network in the same way as for the attacker.
-Weights are assigned to all nodes based on the attacker strategies, with each attacker strategy given equal probability.
-We develop a novel approximation algorithm for the defender to generate a near-optimal defender strategy.
-Thus, for the defender, the problem can be considered as finding a near-optimal path with the aim of covering at least one node from each attacker strategy.
-For this, we use different colors to represent each attacker strategy.
-The developed defender strategy includes as many different colored vertices as possible.
-We consider a Stackelberg game instead of a zero-sum game.
-The advantage of the Stackelberg game over the zero-sum game is that it allows for a more realistic representation of real-world scenarios where one player has more information, resources, or power than the other player.
-In this Stackelberg game, the defenders have an advantage because they can consider the attacker’s potential moves and responses when making their decisions.
-This can result in a more efficient outcome than in a zero-sum game, where the players are equally matched.
-Due to the complex transportation network and limited police resources, generating efficient strategies for both players is a challenging task.
-We can summarize the main contributions of this paper as follows: (1) To the best of our knowledge, this is the first attempt to solve a Stackelberg game considering the escape interdiction game (EIG) problem using the multi-layer network concept, with a focus on developing solution methodologies.
-(2) We develop an optimal approach for the attacker and a near-optimal approach for the defenders, and we are not aware of any previous implementations for escape interdiction problems on multi-layer network.
-(3) We generate a multi-layer network to create efficient strategies for both players.
-(4) We use grids dataset for the case study and show that our MLN-EIGS approach outperforms MILP-EIGS approach in terms of solution quality, with a minimal performance gap.
-The paper is organized as follows.
-In Section 2, we present the relevant research.
-In Section 3, we define the problem description and modeling.
-In Section 4, we describe the solution methodology.
-Section 5 presents the quality of the results.
-We conclude this research in Section 6.
-2 Related Work
-Security games play an important role in providing social security.
-Zhang et al.
-[3] consider a zero-sum game for the escape interdiction problem and develop a MILP-based solution approach to provide an optimal solution.
-To demonstrate the efficacy of their methodology, they generate optimal solutions on grids of different sizes.
-Samanta et al.
-[1] consider the same zero-sum game problem and develop a meta-heuristic-based solution approach to provide a scalable near-optimal solution in a time-efficient manner.
-For the same escape interdiction problem, Samanta et al.
-[2] develop a simulation-based approach to generate a scalable solution to increase security in a large transportation network.
-Recent literature focuses on Stackelberg games considering the escape interdiction problem to decrease the crime rate in society.
-For example, Zychowski and Mandziuk [4] consider search games (SEG) on directed graphs.
-They consider multiple defender resources and one attacker, where the attacker aims to reach one of several predefined target points from a fixed location.
-They develop a genetic algorithm-based heuristic approach to provide a near-optimal solution on synthetic datasets.
-Basilico et al.
-[9] consider a leader-follower game and formulate the problem as a mathematical programming model.
-They use optimization software tools to solve the model by generating the optimal strategies for both leader and follower.
-Letchford and Conitzer [11] consider security games on graphs and develop a polynomial-time algorithm to generate optimal strategies for players.
-Similarly, Iwashita et al.
-[10] consider security games on graphs and develop an algorithm to reduce the graph by eliminating unnecessary edges, providing a time-efficient, scalable near-optimal solution.
-In the same vein, Karwowski and Mandziuk [18] consider an evader-defender Stackelberg game model and develop a Monte Carlo Tree Search approach to provide efficient patrolling schemes.
-Tsai et al.
-[13] consider an attacker-defender Stackelberg game model and develop a linear program to generate optimal mixed strategies for players by allocating limited resources optimally, with a case study on part of the Mumbai road network.
-Similarly, Shieh et al.
-[12] develop a game-theoretic system to provide security with limited resources in the port of Boston and also test its efficacy in the port of New York.
-They schedule patrols efficiently to provide optimal mixed strategies for players, considering an attacker-defender Stackelberg game model.
-Similar papers focus on Stackelberg security games to provide security in society (e.g., Cermak et al.
-[16], Lou et al.
-[19], Sinha et al.
-[14], ˇCern‘y et al.
-[17], Zhang and Malacaria [15]).
-Conitzer and Sandholm [5] show that exact approaches are not suitable because of the NP-hard nature of this problem, as it faces scalability issues due to the high time complexity.
-Paruchuri et al.
-[6] develop a MILP-based exact algorithm to solve Bayesian Stackelberg security games.
-Bosansky and Cermak [7] also consider Stackelberg games and develop algorithms based on LPs and MILP to generate Strong Stackelberg Equilibrium (SSE) and perform a case study at Los Angeles International Airport, considering security scheduling.
-The layered graph concept is a useful tool for solving complex problems on transportation networks in a time-efficient manner.
-For example, Saito et al.[8] consider the problem of selecting important nodes in a network and construct a layered graph from the original graph, where each layer is added on top as time proceeds to demystify the complex problem.
-Hence, we focus on solving a Stackelberg escape interdiction game using the layered graph concept.
-3 Problem Description
-We consider a two-player Stackelberg game where this sequential game is played between multiple defenders $(D)$ and one attacker $(A)$.
-The defenders have a finite set of actions $\acute{S}$, and the attacker has a finite set of actions represented by $\acute{A}$.
-$S$ and $A$ represent the sets of probability distributions over the defender and attacker actions, respectively.
-The game is modeled as a general-sum game, where $x\in S$ and $y\in A$ are the probability distributions over $\acute{S}$ and $\acute{A}$, respectively.
-For $a_{j}=(v_{j},t^{a}_{j})$ and $s^{r}_{i}=(v^{r}_{i},t^{r,in}_{i},t^{r,out}_{i})$, $d_{r}$ reaches the attacker at node $v^{r}_{i}$ if $v^{r}_{i}$ = $v_{j}$ and $t^{r,in}_{i}\leq t^{a}_{j}\leq t^{r,out}_{i}$.
-The defender gets -1 if the attacker successfully escapes from the transportation network; otherwise, the defender’s utility value is 0 (see Eq.
-1).
-Considering this case, we define $Z_{s^{r}_{i},a_{j}}=1$, otherwise $Z_{s^{r}_{i},a_{j}}=0$.
-Here, $y$ and $A$ are used to calculate the expected utility of the defender, which is represented in Eqs. (2) and (3).
-$U_{d}\;(S,A)=\begin{cases}0,&\text{if $\exists$$Z_{s^{r}_{i},a_{j}}$ = 1, $a_{j}$ $\in$ A, $s^{r}_{i}$ $\in$ $S^{r}$, r $\in$ R}.\\ -1,&\text{ otherwise}.\end{cases}$ (1)
-$U_{d}\;(x,A)=\sum\limits_{S\in\textbf{S}}U_{d}(S,A)\;x_{s}$ (2)
-$U_{d}\;(x,y)=\sum\limits_{A\in\textbf{A}}U_{d}(x,A)\;y_{A}$ (3)
-Therefore, we need an optimization model so that the defenders can find their best optimal strategy by solving the following linear problem (see Eqs. 4-6).
-$max\qquad U^{*}$ (4)
-$s.t.\qquad U^{*}\leq U_{d}\;(x,A)\qquad\forall A\in A^{{}^{\prime}}$ (5)
-$\sum\limits_{S\in\textbf{S}}x_{s}=1,x_{s}\geq 0\qquad\forall S\in S^{{}^{\prime}}$ (6)
-The expected utility of the defender and attacker are $U_{d}$ and $U_{a}$, respectively. For any pair of strategies $(x,y)$, the expected payoff of the defender is denoted by $U_{d}(x,y)$.
-First, in Eq. (7), the defender commits to a strategy $x\in X$, and given such an $x$, the attacker chooses his strategy from the best-response set $BR(x)$ where
-$BR(x)=\;argmax_{y\in Y}\;U_{a}\;(x,y)$ (7)
-Second, in Eq. (8), to maximize utility, the defender chooses the strategy $x$ (the best response of the defender) to the attacker’s best response.
-$max_{x\in S}\;U_{d}\;(x,y)\qquad s.t.\;y\in BR(x)$ (8)
-Third, in a Strong Stackelberg Equilibrium (SSE), we assume that the follower (attacker) breaks ties in favor of the leader (defender) (see Eq. 9). In that case, the optimization problem is
-$max_{x\in S,\;y\in BR(x)}\;U_{d}\;(x,y)$ (9)
-SSE is popular because it is always guaranteed to exist ((Conitzer [20]).
-4 Proposed Mixed Approach on Multi-Layer Networks
-This section presents the proposed methodology for the considered escape interdiction problem which is formulated as a Stackelberg game model.
-A novel mixed approach on layered graph concept is developed to solve the Stackelberg model, where the defender is the leader and the attacker is the follower.
+The defenders attempt to capture the attacker before he can flee the city.
+A novel mixed approach on layered graph concept is developed to solve the Stackelberg model.
 For the defender, the developed approximation algorithm finds the efficient movement of the defender from a random initial location.
@@
 The attacker updates the pure attacker strategy set, where all strategies are of equal probability.
-The algorithm checks if, for $K$ consecutive iterations, the game utility is below a threshold value named $\epsilon$.
-The value of $\epsilon$ is set to 0.05 for our experiments.
-If this convergence criterion is met, the algorithm converges, and we return the best defender strategy with the probability of interdiction and total runtime (see Algorithm 1).
-The flowchart is shown in Fig.
-1.
-Figure 1: Flowchart of Stackelberg Game Model
+The algorithm checks if, for a certain number of consecutive iterations, the game utility is below a threshold value.
+If this convergence criterion is met, the algorithm converges, and we return the best defender strategy with the probability of interdiction and total runtime.
+To set a benchmark, the developed Stackelberg game is solved using the adopted MILP approaches for both the attacker and defender.
+We consider a Stackelberg game instead of a zero-sum game.
+The advantage of the Stackelberg game over the zero-sum game is that it allows for a more realistic representation of real-world scenarios where one player has more information, resources, or power than the other player.
+In this Stackelberg game, the defenders have an advantage because they can consider the attacker’s potential moves and responses when making their decisions.
+This can result in a more efficient outcome than in a zero-sum game, where the players are equally matched.
+Due to the complex transportation network and limited police resources, generating efficient strategies for both players is a challenging task.
+We can summarize the main contributions of this paper as follows: (1) To the best of our knowledge, this is the first attempt to solve a Stackelberg game considering the escape interdiction game (EIG) problem using the multi-layer network concept, with a focus on developing solution methodologies.
+(2) We develop an optimal approach for the attacker and a near-optimal approach for the defenders, and we are not aware of any previous implementations for escape interdiction problems on multi-layer network.
+(3) Using grid datasets, we demonstrate that our method significantly outperforms the MILP approach in both computational time and solution quality, with minimal performance trade-offs.
+The paper is organized as follows.
+In Section 2, we present the relevant research.
+In Section 3, we define the problem description and modeling.
+In Section 4, we describe the solution methodology.
+The benchmarking algorithm is described in Section 5.
+Section 6 presents the quality of the results.
+We conclude this research in Section 7.
+2 Related work
+Security games play an important role in providing social security ([23], [25]).
+Recent literature focuses on Stackelberg games considering the escape interdiction problem to decrease the crime rate in society.
+For example, [4] consider search games (SEG) on directed graphs.
+They consider multiple defender resources and one attacker, where the attacker aims to reach one of several predefined target points from a fixed location.
+They develop a genetic algorithm-based heuristic approach to provide a near-optimal solution on synthetic datasets.
+Similarly, [9] consider a leader-follower game and formulate the problem as a mathematical programming model.
+They use optimization software tools to solve the model by generating the optimal strategies for both leader and follower.
+Again, [11] consider security games on graphs and develop a polynomial-time algorithm to generate optimal strategies for players.
+Similarly, [10] consider security games on graphs and develop an algorithm to reduce the graph by eliminating unnecessary edges, providing a time-efficient, scalable near-optimal solution.
+In the same vein, [18] consider an evader-defender Stackelberg game model and develop a Monte Carlo Tree Search approach to provide efficient patrolling schemes.
+[13] consider an attacker-defender Stackelberg game model and develop a linear program to generate optimal mixed strategies for players by allocating limited resources optimally, with a case study on part of the Mumbai road network.
+Likewise, [12] develop a game-theoretic system to provide security with limited resources in the port of Boston and also test its efficacy in the port of New York.
+They schedule patrols efficiently to provide optimal mixed strategies for players, considering an attacker-defender Stackelberg game model.
+Similar papers focus on Stackelberg security games to provide security in society (e.g., [16], [19], [14], [17], [15]).
+[5] show that exact approaches are not suitable due to the NP-hard nature of this problem, as it faces scalability issues due to the high time complexity.
+On the contrary, [6] develop a MILP-based exact algorithm to solve Bayesian Stackelberg security games.
+Similarly, [7] consider Stackelberg games and develop algorithms based on LPs and MILP to generate Strong Stackelberg Equilibrium (SSE) and perform a case study at Los Angeles International Airport, focusing on security scheduling.
+Considering a zero-sum game for the escape interdiction problem, [3] develop a MILP-based solution approach to provide an optimal solution.
+To demonstrate the efficacy of their methodology, they generate optimal solutions on grids of different sizes.
+[1] consider the same zero-sum game problem and develop a meta-heuristic-based solution approach to provide a scalable near-optimal solution in a time-efficient manner.
+Again, [2] develop a simulation-based approach to generate a scalable solution to increase security in a large transportation network for this escape interdiction problem.
+The layered graph concept is a useful tool for solving complex problems on transportation networks in a time-efficient manner.
+For example, [8] consider the problem of selecting important nodes in a network and construct a layered graph from the original graph, where each layer is added on top as time proceeds to demystify the complex problem.
+We focus on solving a Stackelberg escape interdiction game using the layered graph concept.
+3 Problem description
+We consider a two-player Stackelberg game where this sequential game is played between multiple defenders $(D)=\{d_{r}\mid r\in R\}$ and one attacker $(A)$.
+The defenders have a finite set of actions $\acute{S}$, and the attacker has a finite set of actions represented by $\acute{A}$.
+$S$ and $A$ represent the sets of probability distributions over the defender and attacker actions, respectively.
+The game is modeled as a general-sum game, where $x\in S$ and $y\in A$ are the probability distributions over $\acute{S}$ and $\acute{A}$, respectively.
+The game begins at time 0 and ends at time $t_{max}>0$.
+$v_{\infty}$ signifies any exit node in the considered network.
+The sequence of states $\textit{A}=<a_{1}=(v^{a}_{0},0),...,a_{j}=(v_{j},t^{a}_{j}),...,a_{k}=(v_{\infty},t^{a}_{k}\leq t_{max})>$ signifies the attacker’s pure strategy.
+Each state $a_{j}=(v_{j},t^{a}_{j})$ indicates that at time $t^{a}_{j}$, the attacker is present at node $v_{j}$.
+Likewise, a defender’s state $d_{r}$ is a tuple $S^{r}=(v^{r},t^{r,in},t^{r,out})$, representing the state where defender $d_{r}$ is present at the node $v^{r}$ during the interval $[t^{r,in},t^{r,out}]$.
+The defender’s pure strategy consists of a set of $m$ schedules corresponding to the total number of defenders, $m$.
+The schedule for defender $d_{r}$ is defined as a sequence of states $S^{r}=<s^{r}_{1},...,s^{r}_{i},...,s^{r}_{k}>$, where $s^{r}_{1}=(v^{r}_{0},0,t^{r,out}_{1})$ and $s^{r}_{k}=(v^{r}_{k},t^{r,in}_{k},t_{max})$.
+The mixed strategy for the defender is denoted by $x=<x_{S}>$, where $x_{S}$ represents the probability with which the strategy $S$ is employed.
+For $a_{j}=(v_{j},t^{a}_{j})$ and $s^{r}_{i}=(v^{r}_{i},t^{r,in}_{i},t^{r,out}_{i})$, the defender $d_{r}$ intercepts the attacker at node $v^{r}_{i}$ if $v^{r}_{i}$ = $v_{j}$ and $t^{r,in}_{i}\leq t^{a}_{j}\leq t^{r,out}_{i}$.
+In this context, we define $z_{s^{r}_{i},a_{j}}=1$ if the conditions are met, otherwise $z_{s^{r}_{i},a_{j}}=0$.
+The defender incurs a utility of -1 if the attacker successfully escapes from the transportation network; otherwise, the defender’s utility is 0 (see Eq.
+1).
+Here, $y$ and $A$ are used to calculate the expected utility of the defender, which is represented in Eqs. (2) and (3).
+$U_{d}\;(S,A)=\begin{cases}0,&\text{if $\exists$$z_{s^{r}_{i},a_{j}}$ = 1, $a_{j}$ $\in$ {A}, $s^{r}_{i}$ $\in$ $S^{r}$, r $\in$ R}.\\ -1,&\text{ otherwise}.\end{cases}$ (1)
+$U_{d}\;(x,A)=\sum\limits_{S\in\textbf{S}}U_{d}(S,A)\;x_{s}$ (2)
+$U_{d}\;(x,y)=\sum\limits_{A\in\textbf{A}}U_{d}(x,A)\;y_{A}$ (3)
+Therefore, we need an optimization model so that the defenders can find their best optimal strategy by solving the following linear problem (see Eqs. 4-6).
+$max\qquad U^{*}$ (4)
+$s.t.\qquad U^{*}\leq U_{d}\;(x,A)\qquad\forall A\in A^{{}^{\prime}}$ (5)
+$\sum\limits_{S\in\textbf{S}}x_{s}=1,x_{s}\geq 0\qquad\forall S\in S^{{}^{\prime}}$ (6)
+The expected utility of the defender and attacker are $U_{d}$ and $U_{a}$, respectively (see Eqs.
+7-8).
+For any pair of strategies $(x,y)$, the expected payoff of the defender is denoted by $U_{d}(x,y)$.
+First, considering that the defender commits to a strategy $x\in X$, and given such an $x$, the attacker chooses his strategy from the best-response set $BR(x)$ where
+$BR(x)=\;argmax_{y\in Y}\;U_{a}\;(x,y)$ (7)
+Second, to maximize utility, the defender chooses the strategy $x$ (the best response of the defender) to the attacker’s best response.
+$max_{x\in S}\;U_{d}\;(x,y)\qquad s.t.\;y\in BR(x)$ (8)
+Third, in a SSE, we assume that the follower (attacker) breaks ties in favor of the leader (defender) (see Eq. 9). In that case, the optimization problem is
+$max_{x\in S,\;y\in BR(x)}\;U_{d}\;(x,y)$ (9)
+SSE is popular because it is always guaranteed to exist ([21]). The flowchart of the Stackelberg game model is shown in Fig. 1.
+Figure 1: Flowchart of Stackelberg game model.
+4 Proposed mixed approach on multi-layer networks
+This section presents the proposed methodology for the considered escape interdiction problem which is formulated as a Stackelberg game model.
+A novel mixed approach on layered graph concept is developed to solve the Stackelberg model, where the defender is the leader and the attacker is the follower.
+In this Stackelberg game, first, the defenders commit to a strategy, and then the attacker generates the best response to the given defender strategy.
+We use the concept of a multi-layer network (MLN) in which, at each time-stamp, we create a copy of the entire network.
+We consider edge length as the time factor to create connections between these multi-layer networks.
+This means that depending on the edge length, we choose the layers from which the start and end nodes of that particular edge are selected.
+In this way, for all edges in the original transportation network, we create corresponding edges in the multi-layer network.
+We consider a set of strategies for both players, i.e., the defenders and the attacker.
+For the defender, we use a mixed strategy set where each strategy is assigned a mixed probability, with the sum of these probabilities equaling one.
+For the attacker, we consider a pure strategy set.
+To generate the optimal attacker strategy, we first assign weights to the nodes depending on the mixed probabilities and the nodes’ presence in the defender strategies.
+We check the presence of each node in all defender strategies.
+If the node is present in multiple strategies, we sum the mixed probabilities of those strategies.
+This final probability is then assigned as the weight of the node.
+Then, we assign the weight of each node to all of its incoming edges.
+Finally, we apply Dijkstra’s algorithm on this network to find the optimal strategy for the attacker.
+For the defender, we create a multi-layer network in the same way as for the attacker.
+Weights are assigned to all nodes based on the attacker strategies, with each attacker strategy given equal probability.
+We develop a novel approximation algorithm for the defender to generate a near-optimal defender strategy.
+Thus, for the defender, the problem can be considered as finding a near-optimal path with the aim of covering at least one node from each attacker strategy.
+For this, we use different colors to represent each attacker strategy.
+The developed defender strategy includes as many different colored vertices as possible.
+The algorithm converges if for $K$ consecutive iterations, the game utility is below a threshold value named $\epsilon$.
+Then the best defender strategy with the maximum probability of interdiction and total runtime are returned (see Algorithm 1).
 Input: Initialize the initial strategy sets of defender and attacker $S^{\prime},A^{\prime}$;
@@
 return $U_{d}$.
-Algorithm 1 MLN-EIGS Algorithm
-To create a benchmark, we construct a novel Stackelberg game where the defender is the leader, and the attacker is the follower.
-The best strategies for the attacker and defender are generated using bestAO and bestDo, which are the MILP approaches for the attacker and defender, respectively (see Algorithm 2).
-Zhang et al.
-[3] develop these optimal approaches to provide optimal strategies for attackers and defenders given a strategy set for each player.
-Similar to the MLN-EIGS, in this MILP-EIGS the defender commits to the best mixed strategy.
-The attacker takes this best defender mixed strategy as input, generates the optimal attacker strategy, and updates the pure attacker strategy set.
-Then, we calculate the game utility, and the game continues until the game utility is below $\epsilon$ for $K$ consecutive iterations.
-Here $\epsilon$ acts as a threshold with a value of 0.05.
-We return the defender utility value once the convergence is reached.
-Input: Initialize the initial strategy sets of defender and attacker $S^{\prime},A^{\prime}$;
-Output: $U_{d}:$ Defenders’ game utility;
-repeat
-       $(x,y)\leftarrow CoreLP(S^{\prime},A^{\prime})$;
-       Defender commits to the best strategy $x^{\star}\in x$ which provides maximum utility of defender ;
-       $BR:\;A^{*}\leftarrow{BestAO(x^{\star})}$;
-       if $A^{*}\neq\emptyset$ then
-             $A{{}^{\prime}}\leftarrow A^{{}^{\prime}}\cup A^{*}$;
-       end if
-      $S^{*}\leftarrow{BestDO(A^{{}^{\prime}})}$;
-       if $S^{*}\neq\emptyset$ then
-             $S^{{}^{\prime}}\leftarrow S^{{}^{\prime}}\cup S^{*}$;
-       end if
-until convergence;
-return $U_{d}$.
-Algorithm 2 MILP-EIGS Benchmarking Algorithm
+Algorithm 1 MLN-EIGS algorithm.
 4.1 Efficient attacker strategy design using exact approach on time expanded network
@@
 Update the minimum Prob. of. Interdiction $(1-P)$ value to the node depending on its presence in the defender strategy:
+Algorithm 2 Optimal attacker strategy design using time expanded network.
 Initial probability of interdiction $(P)$ of all nodes in all layered graphs = 0 ;
@@
 end for
-Algorithm 3 Optimal Attacker Strategy Design Using Time Expanded Network
 Update the minimum Prob. of. Interdiction $(1-P)$ value to the node depending on its presence in the defender strategy:
@@
       A threshold value of the path length $t$.
-    Here, each colored vertex corresponds to a particular attacker’s strategy (a path).
+    Here, each colored vertex corresponds to a particular attacker’s strategy.
   • 2.
     Goal:
-    To find a path whose length is at most $t$ and contains as many different colored vertices as possible.
+    To find a defender’s strategy whose length is at most $t$ and contains as many different colored vertices as possible.
   • 3.
     Guess:
-    This problem is difficult as Dijkstra cannot solve its decision version, which is $NP$-complete. We provide a formal proof in this paper.
-We can use the steps below for the approximate algorithm.
+    This problem is difficult (say, its decision version is $NP$-complete) and thus cannot be solved by Dijkstra. We provide a formal proof in this paper.
+We use the steps below for the approximate algorithm.
   • 1.
-    Step1: Find a shortest path to any one of colored vertices from the start vertex (we can use Dijkstra). Assume the path is to vertex $v_{1}$ with color $c_{i_{1}}$.
+    Step1: Find a shortest path to any one of colored vertices from the start vertex using Dijkstra. Assume the path is to vertex $v_{1}$ with color $c_{i_{1}}$.
   • 2.
@@
 If we cannot find a path that covers all colors, we can create a tree (or a set of paths) instead, which corresponds to a non-deterministic strategy for a defender.
-In a simplest form, if the path we obtained in the previous method does not cover a subset of colors, we create another path from the initial vertex, which tries to cover these remaining colors only.
+In the simplest form, if the path we obtained in the previous method does not cover a subset of colors, we create another path from the initial vertex, which tries to cover these remaining colors only.
 We repeat this procedure until all colors are covered.
 Then, the defender flips a coin and chooses one path.
-4.3 Theorem: Problem ”color covering” is $NP$-complete
-Problem ”color covering”
+4.3 Problem ‘color covering’ is $NP$-complete
+Problem ‘color covering’
   • 1.
@@
   • 2.
-    Output: ”yes” when there exists a path from the initial vertex with length $n$, such that all $m$ colors appear at least one of a vertex on the path. ”no” otherwise.
-Theorem: Problem ”color covering” is NP-complete
+    Output: ‘Yes’ if there exists a path from the initial vertex with length $n$, such that all $m$ colors appear on at least one vertex along the path. ‘No’ otherwise.
+Theorem: Problem ‘color covering’ is NP-complete
   • 1.
@@
   • 2.
-    Output: ”yes” if there exists an assignment of variables that makes all clauses true. ”no” otherwise.
+    Output: ‘Yes’ if there exists an assignment of variables that makes all clauses true. ‘No’ otherwise.
 For a given $3-SAT$ instance, we create an $n+1$ level network.
@@
   • 5.
-    If a clause with color $c$ contains $x_{i}$, the ”true” vertex for $x_{i}$ has color $c$.
+    If a clause with color $c$ contains $x_{i}$, the ‘true’ vertex for $x_{i}$ has color $c$.
   • 6.
-    If a clause with color $c$ does not contain $x_{i}$, the ”false” vertex for $x_{i}$ has color $c$.
+    If a clause with color $c$ does not contain $x_{i}$, the ‘false’ vertex for $x_{i}$ has color $c$.
 Thus, the $3-SAT$ instance is satisfiable iff there exists a path from the initial vertex with length $n$, which covers all colors.
+5 MILP-EIGS benchmarking algorithm
+To establish a benchmark, we formulate a novel Stackelberg game in which the defender acts as the leader and the attacker as the follower.
+The optimal strategies for both the attacker and defender are derived using bestAo and bestDo, which represent the MILP approaches for the attacker and defender, respectively (see Algorithm 3).
+These optimal approaches, developed by [3], provide the best strategies for attackers and defenders, given a predefined strategy set for each player.
+Similar to the MLN-EIGS approach, in this MILP-EIGS framework, the defender commits to the optimal mixed strategy.
+The attacker then uses the best defender mixed strategy as input to generate the optimal attacker strategy and updates his pure strategy set accordingly.
+The game utility is computed, and the process continues until the game utility remains below a threshold value $\epsilon$ for $K$ consecutive iterations.
+Once the game converges, the defender’s utility value is returned.
+Input: Initialize the initial strategy sets of defender and attacker $S^{\prime},A^{\prime}$;
+Output: $U_{d}:$ Defenders’ game utility;
+repeat
+       $(x,y)\leftarrow CoreLP(S^{\prime},A^{\prime})$;
+       Defender commits to the best strategy $x^{\star}\in x$ which provides maximum utility of defender ;
+       $BR:\;A^{*}\leftarrow{bestAo(x^{\star})}$;
+       if $A^{*}\neq\emptyset$ then
+             $A{{}^{\prime}}\leftarrow A^{{}^{\prime}}\cup A^{*}$;
+       end if
+      $S^{*}\leftarrow{bestDo(A^{{}^{\prime}})}$;
+       if $S^{*}\neq\emptyset$ then
+             $S^{{}^{\prime}}\leftarrow S^{{}^{\prime}}\cup S^{*}$;
+       end if
+until convergence;
+return $U_{d}$.
+Algorithm 3 MILP-EIGS benchmarking algorithm.
+The vehicle interdiction problem is proved to be NP-hard ([3]).
+The best oracles, that is, the MILPs, encounter significant space and time complexity when applied to moderately large urban road networks with many nodes and edges.
+The MILP consists of bestDo for defenders and bestAo for the attacker.
+[3] develop these MILP approaches to find the best strategies.
+5.1 MILP for attacker: bestAo
+The MILP approach for the attacker (bestAo) constructs an optimal path for the attacker from the crime node to the exit node.
+The attacker’s utility decreases when more defender paths interdict this new attacker path.
+In this formulation, $A_{j,v}$ represents whether the attacker reaches node $v$ in the $j^{th}$ state.
+Equation (11) specifies that the attacker starts at node $v^{a}_{0}$, aims to reach the exit node $v_{\infty}$, and occupies only one node in each state.
+Equations (12) and (13) ensure that once the attacker reaches the sink node $v_{\infty}$, he remains there in subsequent states and confirms that the attacker’s strategy forms a valid path.
+Specifically, the attacker can only reach a neighboring node $u$ in state $j+1$ if they were at node $v$ in state $j$, as indicated by $\omega_{j,(v,u)}$ in Equations (14) and (15).
+Equations (16) and (17) initialize the time and update it in subsequent states.
+Finally, Equations (18) to (20) determine whether the attacker is intercepted by defender paths $S$, denoted by $z_{S}$.
+$max\;\sum\limits_{s\in S^{{}^{\prime}}}(1-z_{s})x_{s}$ (10)
+$s.t.\qquad A_{1,v^{a}_{0}}=1,A_{L^{a}_{max},v_{\infty}}=1,\sum\limits_{v\in V}A_{j,v}=1\qquad\forall j$ (11)
+$A_{j+1,v_{\infty}}\geq A_{j,v_{\infty}}\qquad\forall j$ (12)
+$\sum\limits_{u\in N(v)}A_{j+1,u}\geq A_{j,v}\qquad\forall v\in V,j$ (13)
+$\omega_{j,(v,u)}\leq min(A_{j,v},A_{j+1,u})\qquad\forall(v,u)\in E,j$ (14)
+$\omega_{j,(v,u)}\geq A_{j,v}+A_{j+1,u}-1\qquad\forall(v,u)\in E,j$ (15)
+$t^{a}_{1}=0$ (16)
+$t^{a}_{j+1}\geq t^{a}_{j}+\sum\limits_{(v,u)\in E}t_{(v,u)}\;\omega_{j,(v,u)}\qquad\forall j$ (17)
+$-M\alpha^{j}_{S,r,i}\leq t^{r,in}_{i}-t^{a}_{j}\leq M(1-\alpha^{j}_{S,r,i})\qquad\forall S,r,i,j$ (18)
+$-M\beta^{j}_{S,r,i}\leq t^{a}_{j}-t^{r,out}_{i}\leq M(1-\beta^{j}_{S,r,i})\qquad\forall S,r,i,j$ (19)
+$z_{S}\geq\alpha^{j}_{S,r,i}+\beta^{j}_{S,r,i}+A_{j,v^{S^{r}}_{i}}-2\qquad\forall S,r,i,j$ (20)
+$A_{j,v},\alpha^{j}_{S,r,i},\beta^{j}_{S,r,i},\omega_{j,(v,u)},z_{S}\in\{0,1\},t^{a}_{j}\in[0,t_{max}]$ (21)
+5.2 MILP for defender: bestDo
+The MILP approach for defenders (bestDo) is described below, focusing on maximizing the rate at which the defender can intercept the attacker within a specified time frame in a large transportation network.
+The bestDo provides the optimal formulation for the defender’s movements over time.
+It devises a path for the defender that intercepts the maximum number of attacker paths, thereby maximizing the defender’s utility.
+In this approach, the attacker attempts to flee the city by utilizing any exit node $v_{\infty}$ from the crime scene, represented as node $v^{0}_{a}\in V$.
+A defender’s state is defined by the tuple $(v^{r},t^{r,in},t^{r,out})$, indicating the defender’s presence at node $v^{r}$ during the time interval $[t^{r,in},t^{r,out}]$.
+The waiting time at a node is calculated as $t^{r,out}-t^{r,in}=k\delta$, and the travel time between consecutive nodes is estimated as $t^{r,in}_{i+1}-t^{r,out}_{i}=dist(v^{r}_{i},v^{r}{i+1})/sp_{d_{r}}$.
+In this context, $s^{r}_{i,v}=1$ signifies that defender $d_{r}$ reaches node $v$ at the $i^{th}$ state of $S^{r}$.
+Equation (23) specifies that the initial position of $d_{r}$ is $v^{r}_{0}$, and $d_{r}$ remains on a single node at each state of $S^{r}$.
+In Equations (24) and (25), the $0/1$ variable $\omega_{r,i,(v,u)}$ indicates whether a path exists between nodes $v$ and $u$ between the $i^{th}$ and $(i+1)^{th}$ states of $S^{r}$.
+Equation (26) specifies that the strategy begins at time 0 and ends at $t_{max}$, where $L_{max}$ represents the maximum allowable length of the defender’s strategy sequence, and the time the defender remains at the $i^{th}$ state is $k_{r,i}\delta$.
+Equation (27) guarantees that the defender travels between stops along the shortest path.
+With Eqs.(28)-(32), $z_{A}$ indicates whether the attacker using path $A$ is intercepted.
+Specifically, $(v^{A}_{j},t^{A}_{j})$ represents the attacker’s $j^{th}$ state in $A$, while $\gamma^{A,j}_{r,i}$ indicates whether the attacker using path $A$ encounters defender $d_{r}$ at the attacker’s $j^{th}$ position while $d_{r}$ is at their $i^{th}$ stop.
+Additionally, $\alpha^{A,j}_{r,i}$ and $\beta^{A,j}_{r,i}$ denote whether the attacker arrives at their $j^{th}$ position after $d_{r}$ arrives or before $d_{r}$ leaves, respectively.
+The constant $M$ is a large number used in the formulation.
+$max\;-\sum\limits_{A\in A^{{}^{\prime}}}(1-z_{A})y_{A}$ (22)
+$s.t.\qquad s^{r}_{1,v^{r}_{0}}=1,\sum\limits_{v\in V\setminus\{v_{\infty}\}}s^{r}_{i,v}=1\qquad\forall r,i$ (23)
+$\omega_{r,i,(v,u)}\leq min(s^{r}_{i,v},s^{r}_{i+1,u})\qquad\forall r,i,u,v$ (24)
+$\omega_{r,i,(v,u)}\geq s^{r}_{i,v}+s^{r}_{i+1,u}-1\qquad\forall r,i,u,v$ (25)
+$t^{r,in}_{1}=0,t^{r,out}_{L^{d}_{max}}=t_{max},t^{r,out}_{i}=t^{r,in}_{i}+k_{r,i}\delta\qquad\forall r,i$ (26)
+$t^{r,in}_{i+1}=t^{r,out}_{i}+\sum\limits_{v,u\in V\setminus\{v_{\infty}\}}dist(v,u)\omega_{r,i,(v,u)}\qquad\forall r,i$ (27)
+$-M\alpha^{A,j}_{r,i}\leq t^{r,in}_{i}-t^{A}_{j}\leq M(1-\alpha^{A,j}_{r,i})\qquad\forall r,i,A,j$ (28)
+$-M\beta^{A,j}_{r,i}\leq t^{A}_{j}-t^{r,out}_{i}\leq M(1-\beta^{A,j}_{r,i})\qquad\forall r,i,A,j$ (29)
+$\gamma^{A,j}_{r,i}\leq(\alpha^{A,j}_{r,i}+\beta^{A,j}_{r,i}+s^{r}_{i,v^{A}_{j}})/3\qquad\forall r,i,A,j$ (30)
+$\gamma^{A,j}_{r,i}\geq\alpha^{A,j}_{r,i}+\beta^{A,j}_{r,i}+s^{r}_{i,v^{A}_{j}}-2\qquad\forall r,i,A,j$ (31)
+$z_{A}\leq\sum\limits_{j,r,i}\gamma^{A,j}_{r,i}\qquad\forall A$ (32)
+$s^{r}_{i,v},\omega_{r,i,(v,u)},\alpha^{A,j}_{r,i},\beta^{A,j}_{r,i},\gamma^{A,j}_{r,i},z_{A}\in\{0,1\}$ (33)
+$k_{r,i}\in Z_{\geq 0},t^{r,in}_{i},t^{r,out}_{i}\in[0,t_{max}]$ (34)
+To assess optimality, we compare the utility and computational time of the developed MLN-EIGS with that of MILP-EIGS.
 6 Results and discussion
@@
 In Fig.
-2, we consider a sample network and use Dijkstra’s algorithm on the time-expanded network to generate the optimal attacker strategy.
-Here, we provide two mixed defender strategies as input with probabilities of 0.6 and 0.4.
+2, we consider a sample network of 4 nodes in which police stations are nodes 2 and 3, the crime node is 1, the maximum time limit ($t_{max}$) is 5, and the exit point is node 4.
+Here, $0\_4$ indicates node 4 at timestamp 0 $(t=0)$ in the multi-layer network.
+In this example, we provide two mixed defender strategies as input with probabilities of $1/3$ and $2/3$.
+Each node in this multi-layer network is assigned a corresponding probability of interdiction.
+To generate the optimal attacker strategy, we use Dijkstra’s algorithm on the time-expanded network (see Fig.
+3).
+The final attacker strategy is represented by the red line in the multi-layer network, which follows the path $0\_1\rightarrow 3\_3\rightarrow 5\_4$.
 We demonstrate that our developed exact approach for the attacker can generate the optimal attacker strategy, enabling the attacker to escape without interdiction in a concise amount of time (see Table 1).
-Figure 2: Optimal Attacker Strategy Design Using Dijkstra Algorithm on Time Expanded Network.
+Figure 2: Sample network for designing the optimal attacker strategy.
+Figure 3: Design of a multi-layer network for attacker considering a sample network (Fig. 2).
 Table 1:
@@
 Game Parameters
-Network Size: Crime Node: $0\_1$, Police Station: $0\_6$, $T_{max}$: 6, Exit Point: 5
+Network Size: 4 Nodes, Crime Node: $0\_1$, Police Stations: $0\_2$, $0\_3$, $T_{max}$: 5, Exit Point: 4
 Test Case Optimal Attacker strategy Utility of the Final Optimal Attacker Strategy Run Time (Sec)
@@
 In Fig.
-3, we consider a sample network and use the approximation algorithm on the time-expanded network to generate the near-optimal defender strategy.
-Here, we provide three pure attacker strategies as input with equal probability.
+4, we consider a sample network of 6 nodes in which the police station is node 6, the crime node is 1, the maximum time limit $t_{max}$ is 6, and the exit point is node 5.
+Here, $0\_6$ indicates node 6 at timestamp 0 $(t=0)$ in the multi-layer network.
+We input three attacker strategies, each with an equal probability.
+Each node within the same attacker strategy is colored identically in this multi-layer network.
+We use an approximation algorithm on the time-expanded network to generate the near-optimal defender strategy (see Fig.
+5).
+The final defender strategy is represented by the green curvy lines in the multi-layer network, which follows the path $0\_6\rightarrow 2\_3\rightarrow 4\_4\rightarrow 5\_5\rightarrow 6\_5$ and $0\_6\rightarrow 2\_3\rightarrow 4\_5\rightarrow 5\_5\rightarrow 6\_5$.
 We demonstrate that our developed approach for the defender can generate an efficient defender strategy that interdicts all attacker strategies quickly (see Table 2).
-Figure 3: Defender Strategy Design Using Approximation algorithm.
+Figure 4: Sample network for designing near-optimal defender strategy.
 Table 2:
@@
 Game Parameters
-Network Size: Crime Node: $0\_1$, Police Station: $0\_6$, $T_{max}$: 6, Exit Point: 5
-Test Case Final Defender strategy No. of Non-interdicted Attacker Strategies by the Final Defender Strategy Run Time (Sec)
+Network Size: 6 Nodes, Crime Node: $0\_1$, Police Station: $0\_6$, $T_{max}$: 6, Exit Point: 5
+Test Case Final Defender strategy Utility of the Final Defender Strategy Run Time (Sec)
 1 [$0\_6,2\_3,4\_4,5\_5,6\_5$] 0.0 0.0086
@@
 10 [$0\_6,2\_3,4\_5,5\_5,6\_5$] 0.0 0.0081
-We compare the final defenders’ utility generated by the developed MLN-EIGS algorithm with the final defenders’ utility using MILP-EIGS algorithm, which uses the exact approaches named bestDo and bestAo, developed by Zhang et al.
-[3].
-To establish a benchmark, we solve the Stackelberg Game using the bestDo and bestAo exact approaches.
+Figure 5: Design of a multi-layer network for defender considering a sample network (Fig. 4).
+We compare the final defenders’ utility generated by the developed MLN-EIGS algorithm with the final defenders’ utility using the MILP-EIGS algorithm, which employs the exact approaches named bestDo and bestAo, developed by [3].
+To establish a benchmark, we solve the Stackelberg Game using the bestDo and bestAo MILP approaches.
+In both cases, the Stackelberg Game converges after $K$ consecutive iterations where the game utility is below $\epsilon$.
+Here, the values of $K$ and $\epsilon$ are assigned to 10 and 0.05, respectively.
 We compare both the solution quality and computational time.
@@
 10 0.0 0.0 577.01 81.17
-We show that for these small-size grids, our developed mixed approach named MLN-EIGS can provide an optimal solution like the MILP approach named MILP-EIGS in most cases and takes less computational time compared to MILP-EIGS.
-6 Conclusion and Future Directions
-In this paper, we consider a Stackelberg game theoretic model to solve security games considering the dynamic crime scenario.
-This paper proposes a novel optimal routing and resource allocation approach on a transportation network.
-The optimal strategy for the attacker is generated using an exact approach.
-To the best of our knowledge, this is the first time an exact approach using the multi-layer network concept is used for the attacker to solve a security game.
-This developed exact approach consists of Dijkstra’s algorithm and the layered graph concept.
-The near-optimal strategy for the defender is generated using a novel approximation algorithm and the layered graph concept on a transportation network in a time-efficient manner.
-The experiments are conducted on 3x3 to 9x9 grids.
-The results are compared with the solutions developed using MILP approaches for both attacker and defender in this Stackelberg game scenario.
-For benchmarking, we formulate the Stackelberg game to provide optimal solutions using the adopted bestAO and bestDO approaches.
-The contribution of the paper is that the developed MLN-EIGS can offer quality solutions like MILP-EIGS but in a significantly shorter amount of time.
-The developed approaches have a few limitations.
-For example, the proposed approach does not include real-time traffic.
-Future research can focus on developing efficient algorithms to incorporate traffic into the developed mixed approach.
-In addition, an exact approach for the defender can also be developed to generate an optimal strategy for the defender on a time-layered network.
+In the above tables, we consider seven grids ranging in size from 3*3 to 9*9 and evaluate the utility and runtime for the MLN-EIGS and MILP-EIGS approaches by varying the crime node location, initial defender positions, $t_{max}$ (maximum runtime), and exit node.
+In the case of a 3*3 grid, MLN-EIGS takes slightly less computational time with a performance gap of zero, whereas for the 4*4 grid, MILP-EIGS takes nearly 10 seconds longer than MLN-EIGS with no performance gap.
+It is evident that MLN-EIGS provides optimal results similar to MILP-EIGS for smaller grids.
+In the case of the 5*5 grid, MILP-EIGS takes more than 20 seconds longer, and for the 6*6 grid, the time difference exceeds 30 seconds, sometimes reaching 70 seconds.
+MLN-EIGS performs as well as MILP-EIGS for both the 5*5 and 6*6 grids, with only one test case showing a performance gap.
+For the 7*7 and 8*8 grids, while the performance gap remains zero, there is a significant difference in computational time between the two approaches.
+On the 7*7 grid, MILP-EIGS takes over one minute, and on the 8*8 grid, it takes over five minutes longer than MLN-EIGS.
+For the 9*9 grid, although test case four shows a performance gap of 0.5, MLN-EIGS still outperforms MILP-EIGS in terms of computational efficiency.
+Overall, the results indicate that MLN-EIGS consistently provides a time-efficient, high-quality solution, while MILP-EIGS struggles to offer a time-efficient solution for the developed Stackelberg game.
+7 Conclusion
+In this paper, we present a Stackelberg game model to address security challenges in a dynamic crime scenario.
+We propose a novel optimal routing and resource allocation approach on a transportation network.
+The attacker’s optimal strategy is determined using an exact method, which, to our knowledge, is the first to apply the multi-layer network concept in solving a security game.
+This exact approach combines Dijkstra’s algorithm with the layered graph model.
+The defenders’ near-optimal strategy is generated through a novel approximation algorithm, utilizing the layered graph concept on a transportation network in a time-efficient way.
+We solve the developed Stackelberg game using the adopted MILP approaches to provide optimal solutions.
+For benchmarking, we compare the results of our approach with those generated using MILP methods.
+The key contribution of this work is that our MLN-EIGS approach delivers high-quality solutions comparable to MILP-EIGS but in significantly less time.
+There are some limitations, such as the exclusion of real-time traffic.
+Future research could focus on developing efficient algorithms to incorporate traffic data and on creating an exact approach for the defender to generate optimal strategies using the time-layered network concept.
+Acknowledgements We are grateful to the members of the Multi-Agent Laboratory at Kyushu University for their insightful discussions and comments.
+This research is funded by a project supported by the Grants-in-Aid for Scientific Research from the Japan Society for the Promotion of Science.

### v3

@@
 A novel mixed approach on layered graph concept is developed to solve the Stackelberg model.
+This mixed approach employs an exact method for generating the attacker’s strategy and an approximation method for determining the defenders’ strategies.
 For the defender, the developed approximation algorithm finds the efficient movement of the defender from a random initial location.
@@
 Due to the complex transportation network and limited police resources, generating efficient strategies for both players is a challenging task.
-We can summarize the main contributions of this paper as follows: (1) To the best of our knowledge, this is the first attempt to solve a Stackelberg game considering the escape interdiction game (EIG) problem using the multi-layer network concept, with a focus on developing solution methodologies.
-(2) We develop an optimal approach for the attacker and a near-optimal approach for the defenders, and we are not aware of any previous implementations for escape interdiction problems on multi-layer network.
-(3) Using grid datasets, we demonstrate that our method significantly outperforms the MILP approach in both computational time and solution quality, with minimal performance trade-offs.
+This paper introduces a Stackelberg game model for dynamic crime scenarios, presenting an exact attacker strategy based on a multi-layer network and a novel approximation algorithm for the defenders.
+The proposed MLN-EIGS approach produces high-quality solutions comparable to MILP-based methods, while significantly reducing computational time.
 The paper is organized as follows.
@@
 They develop a genetic algorithm-based heuristic approach to provide a near-optimal solution on synthetic datasets.
-Similarly, [9] consider a leader-follower game and formulate the problem as a mathematical programming model.
+[9] consider a leader-follower game and formulate the problem as a mathematical programming model.
 They use optimization software tools to solve the model by generating the optimal strategies for both leader and follower.
@@
 In the same vein, [18] consider an evader-defender Stackelberg game model and develop a Monte Carlo Tree Search approach to provide efficient patrolling schemes.
+In addition, [23] introduce a repeated Stackelberg security game that incorporates a cooperative human behavior model to enhance patrolling strategies for wildlife protection by modeling human decision-making in repeated interactions, thereby improving defender effectiveness against adaptive adversaries in dynamic environments.
 [13] consider an attacker-defender Stackelberg game model and develop a linear program to generate optimal mixed strategies for players by allocating limited resources optimally, with a case study on part of the Mumbai road network.
@@
 We focus on solving a Stackelberg escape interdiction game using the layered graph concept.
-3 Problem description
-We consider a two-player Stackelberg game where this sequential game is played between multiple defenders $(D)=\{d_{r}\mid r\in R\}$ and one attacker $(A)$.
+3 Problem description and modeling
+We consider a two-player Stackelberg game, where the sequential interaction occurs between multiple defenders $\overline{D}=\{d_{r}\mid r\in R\}$ and a single attacker $\overline{A}$.
+The total number of defenders is $m$, and the set of all defenders is represented by $\overline{D}$.
+Here, $r\in R=\{1,\dots,m\}$.
+Since the defenders act jointly and a complete defender strategy comprises the strategies of all individual defenders, while there is only one attacker, we refer to this as a two-player Stackelberg game.
 The defenders have a finite set of actions $\acute{S}$, and the attacker has a finite set of actions represented by $\acute{A}$.
-$S$ and $A$ represent the sets of probability distributions over the defender and attacker actions, respectively.
-The game is modeled as a general-sum game, where $x\in S$ and $y\in A$ are the probability distributions over $\acute{S}$ and $\acute{A}$, respectively.
-The game begins at time 0 and ends at time $t_{max}>0$.
+The game is modeled as a general-sum game, where $x$ and $y$ are the probability distributions over $\acute{S}$ and $\acute{A}$, respectively.
+The transportation network is represented as a directed graph $G=(V,E)$, where $E$ is the set of directed edges corresponding to roads, and $V$ is the set of nodes representing intersections.
+There is a set of predefined exit points in the considered network.
 $v_{\infty}$ signifies any exit node in the considered network.
-The sequence of states $\textit{A}=<a_{1}=(v^{a}_{0},0),...,a_{j}=(v_{j},t^{a}_{j}),...,a_{k}=(v_{\infty},t^{a}_{k}\leq t_{max})>$ signifies the attacker’s pure strategy.
+The game begins at time 0 and ends at time $t_{\text{max}}>0$.
+The sequence of states $A=\langle a_{1}=(v^{a}_{0},0),\ldots,a_{j}=(v_{j},t^{a}_{j}),\ldots,a_{k}=(v_{\infty},t^{a}_{k}\leq t_{\text{max}})\rangle$ represents the pure strategy of the attacker.
 Each state $a_{j}=(v_{j},t^{a}_{j})$ indicates that at time $t^{a}_{j}$, the attacker is present at node $v_{j}$.
 Likewise, a defender’s state $d_{r}$ is a tuple $S^{r}=(v^{r},t^{r,in},t^{r,out})$, representing the state where defender $d_{r}$ is present at the node $v^{r}$ during the interval $[t^{r,in},t^{r,out}]$.
-The defender’s pure strategy consists of a set of $m$ schedules corresponding to the total number of defenders, $m$.
+The defender’s pure strategy is denoted by $S$.
+A pure strategy for the defender consists of $m$ schedules, i.e., $S=\{S^{r}:r\in R\}$.
 The schedule for defender $d_{r}$ is defined as a sequence of states $S^{r}=<s^{r}_{1},...,s^{r}_{i},...,s^{r}_{k}>$, where $s^{r}_{1}=(v^{r}_{0},0,t^{r,out}_{1})$ and $s^{r}_{k}=(v^{r}_{k},t^{r,in}_{k},t_{max})$.
-The mixed strategy for the defender is denoted by $x=<x_{S}>$, where $x_{S}$ represents the probability with which the strategy $S$ is employed.
+The mixed strategy for the defender is denoted by $x=<x_{S}>$, where $x_{S}$ represents the probability with which the strategy $S$ is played.
 For $a_{j}=(v_{j},t^{a}_{j})$ and $s^{r}_{i}=(v^{r}_{i},t^{r,in}_{i},t^{r,out}_{i})$, the defender $d_{r}$ intercepts the attacker at node $v^{r}_{i}$ if $v^{r}_{i}$ = $v_{j}$ and $t^{r,in}_{i}\leq t^{a}_{j}\leq t^{r,out}_{i}$.
@@
 $U_{d}\;(S,A)=\begin{cases}0,&\text{if $\exists$$z_{s^{r}_{i},a_{j}}$ = 1, $a_{j}$ $\in$ {A}, $s^{r}_{i}$ $\in$ $S^{r}$, r $\in$ R}.\\ -1,&\text{ otherwise}.\end{cases}$ (1)
-$U_{d}\;(x,A)=\sum\limits_{S\in\textbf{S}}U_{d}(S,A)\;x_{s}$ (2)
-$U_{d}\;(x,y)=\sum\limits_{A\in\textbf{A}}U_{d}(x,A)\;y_{A}$ (3)
-Therefore, we need an optimization model so that the defenders can find their best optimal strategy by solving the following linear problem (see Eqs. 4-6).
+$U_{d}\;(x,A)=\sum\limits_{S\in\acute{S}}U_{d}(S,A)\;x_{s}$ (2)
+$U_{d}\;(x,y)=\sum\limits_{A\in\acute{A}}U_{d}(x,A)\;y_{A}$ (3)
+Here, we define an optimization model that enables the defenders to determine their optimal strategy by solving the following linear program (see Eqs. 4–6).
 $max\qquad U^{*}$ (4)
-$s.t.\qquad U^{*}\leq U_{d}\;(x,A)\qquad\forall A\in A^{{}^{\prime}}$ (5)
-$\sum\limits_{S\in\textbf{S}}x_{s}=1,x_{s}\geq 0\qquad\forall S\in S^{{}^{\prime}}$ (6)
+$s.t.\qquad U^{*}\leq U_{d}\;(x,A)\qquad\forall A\in\acute{A}$ (5)
+$\sum\limits_{S\in\acute{S}}x_{s}=1,x_{s}\geq 0\qquad\forall S\in\acute{S}$ (6)
 The expected utility of the defender and attacker are $U_{d}$ and $U_{a}$, respectively (see Eqs.
@@
 For any pair of strategies $(x,y)$, the expected payoff of the defender is denoted by $U_{d}(x,y)$.
-First, considering that the defender commits to a strategy $x\in X$, and given such an $x$, the attacker chooses his strategy from the best-response set $BR(x)$ where
-$BR(x)=\;argmax_{y\in Y}\;U_{a}\;(x,y)$ (7)
+First, considering that the defender commits to a strategy $x\in\acute{S}$, and given such an $x$, the attacker chooses his strategy from the best-response set $BR(x)$ where
+$BR(x)=\;argmax_{y\in\acute{A}}\;U_{a}\;(x,y)$ (7)
+Given $x$ and $y$, the expected utility of the attacker is defined as $U_{a}(x,y)$.
 Second, to maximize utility, the defender chooses the strategy $x$ (the best response of the defender) to the attacker’s best response.
-$max_{x\in S}\;U_{d}\;(x,y)\qquad s.t.\;y\in BR(x)$ (8)
-Third, in a SSE, we assume that the follower (attacker) breaks ties in favor of the leader (defender) (see Eq. 9). In that case, the optimization problem is
-$max_{x\in S,\;y\in BR(x)}\;U_{d}\;(x,y)$ (9)
-SSE is popular because it is always guaranteed to exist ([21]). The flowchart of the Stackelberg game model is shown in Fig. 1.
-Figure 1: Flowchart of Stackelberg game model.
+$max_{x\in\acute{S}}\;U_{d}\;(x,y)\qquad s.t.\;y\in BR(x)$ (8)
+Third, in case of a Strong Stackelberg Equilibrium (SSE), we assume that the follower (attacker) breaks ties in favor of the leader (defender) (see Eq. 9). In that case, the optimization problem is
+$max_{x\in\acute{S},\;y\in BR(x)}\;U_{d}\;(x,y)$ (9)
+SSE is popular because it is always guaranteed to exist ([20]).
 4 Proposed mixed approach on multi-layer networks
-This section presents the proposed methodology for the considered escape interdiction problem which is formulated as a Stackelberg game model.
+This section presents the proposed solution methodology for the considered escape interdiction problem which is formulated as a Stackelberg game model.
 A novel mixed approach on layered graph concept is developed to solve the Stackelberg model, where the defender is the leader and the attacker is the follower.
@@
 We check the presence of each node in all defender strategies.
-If the node is present in multiple strategies, we sum the mixed probabilities of those strategies.
-This final probability is then assigned as the weight of the node.
-Then, we assign the weight of each node to all of its incoming edges.
-Finally, we apply Dijkstra’s algorithm on this network to find the optimal strategy for the attacker.
+If the node is present in multiple defender strategies, we sum the mixed probabilities of those strategies.
+This resulting probability $P$ is then assigned as the weight of the node.
+Subsequently, a weight of $P$ is assigned to all of its incoming edges, enabling Dijkstra’s algorithm to identify the attacker’s optimal path with the minimum probability of interdiction.
+CoreLP computes the mixed strategies of the defender and attacker, $x$ and $y$, by solving the linear program defined in Eqs.
+(4)–(6), using the complete strategy spaces of the defender and attacker, denoted by $\acute{S}$ and $\acute{A}$, respectively, as input.
+Here, $ExactAO$ denotes the exact approach developed for the attacker, while $ApproxDO$ denotes the approximation algorithm developed for the defender.
 For the defender, we create a multi-layer network in the same way as for the attacker.
@@
 Algorithm 1 MLN-EIGS algorithm.
+The flowchart of the Stackelberg game model is shown in Fig. 1.
+Figure 1: Flowchart of Stackelberg game model.
 4.1 Efficient attacker strategy design using exact approach on time expanded network
-The optimal attacker strategy is generated using the exact approach for the attacker (see Algorithm 2).
-For each time-stamp, we create a copy of the entire network.
-Then, we establish connections in this multi-layer network using the edge lengths in the original graph.
-This algorithm describes how we assign the probability of interdiction to each node.
-Finally, we apply Dijkstra’s algorithm on this multi-layer network to generate the optimal strategy for the attacker.
+The optimal attacker strategy is computed using an exact approach (see Algorithm 2).
+For each time-stamp, a separate copy of the entire network is created, forming a multi-layer structure.
+Connections between these layers are established based on the edge lengths in the original graph.
+Initially, all edge weights in the multi-layer network are set to zero.
+The algorithm outlines the process of assigning interdiction probabilities to each node and the corresponding weights to its adjacent edges.
+Finally, Dijkstra’s algorithm is applied to this multi-layer network to determine the attacker’s optimal strategy.
 Input: Crime node is the “START” node, all exit nodes are the “GOAL” nodes, Original graph $(G_{0})$;
@@
 for $(i=0;i<t_{max};i++)$ do
-       Generate one copy of the original graph/network, labelled as $G_{i+1}$ ;
+       Generate one copy of the original graph/network, labeled as $G_{i+1}$ ;
 end for
@@
 end for
-Update the minimum Prob. of. Interdiction $(1-P)$ value to the node depending on its presence in the defender strategy:
+Update the maximum probability of interdiction for each node:
 Algorithm 2 Optimal attacker strategy design using time expanded network.
-Initial probability of interdiction $(P)$ of all nodes in all layered graphs = 0 ;
+Initial probability of interdiction $(P)$ of all nodes in the layered graph = 0 ;
+Initial weight of all edges in the layered graph = 0 ;
 for all $N$ defender strategies in the best strategy set do
@@
              $Node=n,t_{in}=In\;Time,t_{out}=Out\;Time$ while $t_{in}<t_{out}$ do
-                   Select node $n$ of graph $G_{t_{i}n}$ Update the minimum Prob.
-of Interdiction of that node $n:(P=P+P_{mix})$ where $P_{mix}$ is the mixed prob.
-of the current defender strategy.
-$t_{in}=t_{in}+1$
+                   Select node $n$ of graph $G_{t_{in}}$;
+                   Update the maximum probability of Interdiction of that node $n:(P=P+P_{mix})$ where $P_{mix}$ is the mixed prob. of the current defender strategy.;
+                   $t_{in}=t_{in}+1$;
              end while
@@
 end for
-Update the minimum Prob. of. Interdiction $(1-P)$ value to the node depending on its presence in the defender strategy:
+Update the weights of all incoming edges to the node with interdiction probability $P$ as $P$:
 for all nodes present in the layered graph do
-       if Prob. of interdiction of a node $n$ is $P$ then
+       if Probability of interdiction of a node $n$ is $P$ then
              Assign the weight $P$ to all in-coming edges of that node.
@@
 Apply Dijkstra’s algorithm on the developed Multi-Layer Network.
-return $U_{d}$.
+return The optimal attacker path with the minimum probability of interdiction.
 4.2 Efficient defender strategy design using approximation algorithm on time expanded network
@@
     Step3: Find a shortest path to any one colored vertices except $c_{i_{1}}$ and $c_{i_{2}}$ from $v_{2}$, and so on, until all colors are visited or the total path length reaches $t$.
-If we cannot find a path that covers all colors, we can create a tree (or a set of paths) instead, which corresponds to a non-deterministic strategy for a defender.
 In the simplest form, if the path we obtained in the previous method does not cover a subset of colors, we create another path from the initial vertex, which tries to cover these remaining colors only.
@@
 $k_{r,i}\in Z_{\geq 0},t^{r,in}_{i},t^{r,out}_{i}\in[0,t_{max}]$ (34)
-To assess optimality, we compare the utility and computational time of the developed MLN-EIGS with that of MILP-EIGS.
+To evaluate the efficacy of our proposed approach, MLN-EIGS, we compare its utility and computational time with those of the adopted method, MILP-EIGS.
 6 Results and discussion
@@
 7 Conclusion
-In this paper, we present a Stackelberg game model to address security challenges in a dynamic crime scenario.
-We propose a novel optimal routing and resource allocation approach on a transportation network.
-The attacker’s optimal strategy is determined using an exact method, which, to our knowledge, is the first to apply the multi-layer network concept in solving a security game.
-This exact approach combines Dijkstra’s algorithm with the layered graph model.
-The defenders’ near-optimal strategy is generated through a novel approximation algorithm, utilizing the layered graph concept on a transportation network in a time-efficient way.
-We solve the developed Stackelberg game using the adopted MILP approaches to provide optimal solutions.
-For benchmarking, we compare the results of our approach with those generated using MILP methods.
-The key contribution of this work is that our MLN-EIGS approach delivers high-quality solutions comparable to MILP-EIGS but in significantly less time.
-There are some limitations, such as the exclusion of real-time traffic.
-Future research could focus on developing efficient algorithms to incorporate traffic data and on creating an exact approach for the defender to generate optimal strategies using the time-layered network concept.
-Acknowledgements We are grateful to the members of the Multi-Agent Laboratory at Kyushu University for their insightful discussions and comments.
+We address a Stackelberg game formulation of the escape interdiction game (EIG) using the multi-layer network concept, focusing on the development of effective solution methodologies.
+An exact approach is proposed for the attacker and a near-optimal approach for the defenders.
+To the best of our knowledge, this is the first implementation of the escape interdiction problem using a multi-layer network framework.
+Our paper presents a proof establishing the NP-completeness of the ’color covering’ problem.
+Using grid datasets, we demonstrate that our method significantly outperforms the MILP approach in both computational time and solution quality, with minimal performance trade-offs.
+Some limitations remain, such as the exclusion of real-time traffic data.
+Future research could aim to develop efficient algorithms that incorporate traffic dynamics and to design an exact approach for the defender to generate optimal strategies using the time-layered network model.
+Acknowledgments We are grateful to the members of the Multi-Agent Laboratory at Kyushu University for their insightful discussions and comments.
 This research is funded by a project supported by the Grants-in-Aid for Scientific Research from the Japan Society for the Promotion of Science.